OpenStudy (anonymous):

Find the area of the rhombus. Leave your answer in simplest radical form. a.50 b.10√3 c.25√6 d.50√3

OpenStudy (anonymous):



OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@sai1234

OpenStudy (anonymous):

@Liltico what exactly is 60 and5?

OpenStudy (anonymous):

60 degrees and 5 is 5

OpenStudy (anonymous):



OpenStudy (anonymous):

is the 5 a diagonal or a side?

OpenStudy (anonymous):

side

OpenStudy (anonymous):

then your area is : \[s ^{2} \sin a\] where s is length of any side and a is any interior angle, can you proceed from there?

OpenStudy (anonymous):

i still don't get the right answer

OpenStudy (anonymous):

remember \[\sin 60 = (\sqrt{3})/2\]